### v1

# Stochastic activations

Abstract
[…]
RELU: 1p1-p [Uncaptioned image]SILU: $p$
Figure 1: Stochastic activation randomly selects one of two activations when $x<0$:
(1) RELU selected with probability 1p1-p; otherwise (2) another activation, in particular SILU.
1 Introduction
Large language models (LLMs) (Devlin et al., 2019; Chowdhery et al., 2022; Brown et al., 2020; Vaswani et al., 2017) have revolutionized natural language processing, enabling unprecedented capabilities in text generation, comprehension, and reasoning. Their success stems from scaling model parameters and leveraging vast amounts of data, but this comes with a significant computational complexity. As the demand for more efficient and powerful models grows, researchers are increasingly focused on optimizing their training processes to balance performance with resource constraints.
The majority of the LLM parameters are in the Feed-Forward Network (FFN) layers, where they memorize the training data. FFNs are two linear layers separated by an activation function, and sometimes an additional linear layer that serves as a gating operation. The activation is a $\mathbb{R}\rightarrow\mathbb{R}$ non-linear function. In this context, the choice of activation function plays a crucial role for both the model’s expressivity and efficiency. The simplest activation is RELU (Rectified Linear Unit), that allows positive values to pass through and forces negative inputs to zero. RELU is sparsity-inducing, since, on average, half of its outputs are zero (in practice significantly more). Within a two-layer Multilayer Perceptron, this means that inference on the second layer is a matrix-sparse vector multiplication, so it can be implemented with fewer FLOPs than a matrix-dense vector multiplication. Note that effectively improving the runtime with this sparsity pattern remains challenging.
In practice, the Sigmoid Linear Unit (SILU) activation, combined with a gated design, has consistently outperformed RELU in terms of model accuracy (Shazeer, 2020). Unfortunately, SILU does not induce sparsity. One plausible explanation for RELU’s underperformance is that its gradient for negative input values is zero, which hinders optimization by preventing weight updates in a significant portion of the network. Solutions like Leaky RELU (Maas et al., 2013) circumvent this problem by ensuring non-zero gradient almost everywhere, but they are inferior to SILU and involve abandoning sparsity. In contrast, if the so-called “dying RELU problem” optimization challenge could be effectively addressed, RELU’s theoretical advantages – such as sparsity and computational efficiency – may translate into performance comparable to SILU for a lower number of FLOPs. This disparity presents a challenge: how to harness the efficiency benefits of sparse activations, such as RELU, without sacrificing the empirical advantages of SILU. This motivates our exploration of alternative training strategies that mitigate RELU’s limitations while preserving its benefits.
In this work, we consider two ways to approach this problem. The first approach is activation fine-tuning, denoted Swi+FT: we pre-train the model with an activation that facilitates efficient large language model optimization, then we change the activation to RELU and adapt the model by fine-tuning it further. Our second approach, referred to as StochA (stochastic activations), is a novel technique that randomly selects between multiple activations, either at train or test time. Both approaches allow models to benefit from the superior optimization properties of SILU. These hybrid strategies combine the best of both worlds – maintaining high model performance while unlocking the computational efficiency of sparse activations.
In summary, this paper makes the following contributions:
  • We introduce and analyze two strategies that employ different activation functions at training and inference time, namely Swi+FT and StochA. Both are complementary and make it possible to use activations at inference time that differ from those employed during pre-training.
  • We produce RELU-based models that are much better than those obtained with regular training, i.e., our methods significantly outperform training with RELU only.
  • We show that stochastic activations, when used at inference time, provides an alternative way to generate diverse sequences compared to traditional temperature sampling or other variants.
2 related work
Standard activation functions
Activations are at the core of deep learning, in that they are needed to depart from defining a linear function with limited expressivity. While early neural network architectures were inspired by logistic regression, such as sigmoidal and tanh activations, many activation functions have been evaluated for the Feed Forward layers (FFNs) of transformers. Vaswani et al. (2017) used RELU (Glorot et al., 2011). However, using the RELU activation function leads to some neurons getting stuck in the negative region. As a consequence, they stop learning entirely, since the gradient is zero for negative inputs, and their weights do not get updated. In contrast, Touvron et al. (2023) used SILU as the activation function for the FFN layers of the transformer for the first Llama models. Shazeer (2020) discusses the benefits of SWIGLU, which consists of SILU with gating. There exist many other activation functions such as the Gaussian Error Linear Unit (GELU) (Hendrycks & Gimpel, 2016), Scaled Exponential Linear Unit (SELU) (Klambauer et al., 2017), Swish (Ramachandran et al., 2018) and gated SILU, among others.
In particular, the Leaky RELU (Maas et al., 2013; Redmon et al., 2015; Ridnik et al., 2021; Guo et al., 2024) tried tackling the dying RELU problem by allowing a small, non-zero gradient when the input is negative in order to keep the neurons active and the gradients flowing, reducing the risk of dead neurons.
Adaptive activation functions
Lee et al. (2022) propose an activation, called Adaptive Swish (ASH), that uses stochastic sampling of the top-k percentile elements. It is a generalization of the Swish (Ramachandran et al., 2018) activation function, that uses adaptive thresholding and selects only the values in the top percentiles and sets these to zero otherwise. The threshold is not fixed but depends on the distribution of the input making it an example of a stochastic activation.
Dropout and structured Dropout variants
In the original dropout paper (Srivastava et al., 2014), the authors propose a regularization technique to reduce overfitting and improve generalization of a neural network. It consists of setting to zero a subset of neurons at each training step. Consequently, the dropped neurons do not contribute to the forward pass or receive weight updates during back-propagation. At inference time, all neurons are used and their outputs are scaled by the dropout probability.
LayerDrop (Fan et al., 2019) randomly drops entire layers during training, hence, it encourages the model to be robust to missing layers. At inference time, some layers can be pruned, trading off between speed and accuracy as needed. While the method does not make the model sparse in the usual sense, it induces structured sparsity in the computation graph during training. Other works also introduce structured dropout variants such as DropBlock (Ghiasi et al., 2018), Bayesian dropout (Gal & Ghahramani, 2016), or Beit (Bao et al., 2021) and masked-autoencoder (He et al., 2022) in computer vision, among others.
Quantization approaches
Fan et al. (2020) propose Quant-noise, that mimicks quantization during training by introducing noise to a random subset of weights for each forward pass enabling high compression ratios while maintaining the original model performance. It uses the Straight-Through estimator (STE) (Bengio et al., 2013; Hinton, 2012) to compute the gradients. This training technique ensures that the model is pretrained to observe both the train-time (unquantized) and the inference-time (quantized) models. This ensures proper optimization, bypassing the flat gradient caused by quantization and reducing the discrepancy that results from the late quantization of the model weights.
Sparsity by design
Some works propose to enable sparsity directly in the architecture, for instance, the Mixture of Experts (MoE) or the Product-Key Memory (PKM). The PKM architecture (Lample et al., 2019) uses a memory layer for neural networks which enables the model to access a large learnable memory and thus, it enables long term memory capabilities. It leverages product quantization (PQ) (Jégou et al., 2011) by splitting the key in two parts and using each part in separate codebooks. The combination of each PQ index enables the model to access a larger memory space efficiently. At each forward pass, only a small subset of the memory is accessed, making it computationally efficient.
Mixture of Expert (MoE) models (Yang et al., 2024; Wei et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024) dynamically select and activate the relevant subset of parameters based on the characteristics of the input data. The MoE approach allows MoE models to expand their capacity without proportionally increasing computational complexity. See Mu & Lin (2025) for an overview of the MoE and references therein.
3 Using different activations at train and test time
This section introduces two strategies for improving the optimization during pretraining using an optimization-compliant activation, while preparing the model to a potentially different activation at test time. First we introduce the Swi+FT fine-tuning approach. Then we introduce our Stochastic Activation StochA.
3.1 Fine-tuning with RELU: Swi+FT
In the following, we use SILU and RELU as our training and inference activations. For reference, they are defined in $\mathbb{R}\rightarrow\mathbb{R}$ as:
$\textrm{{RELU}}(x)=\max(x,0)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \textrm{{SILU}}(x)=x\sigma(x),$ (1)
where $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. We choose these two activations because SILU is one of the best options in terms of accuracy, while RELU is simple and sparse. The two activations are also similar: same asymptotes at -$\infty$ and $\infty$, and the same value at 0. SILU is differentiable twice (unlike RELU) and, interestingly, non-monotonous.
In our proposed approach, the training operates as follows:
  • Most of the training steps (during a proportion $1-\alpha$ of the total number of iterations) are carried out with a first activation that is deemed preferable for training. We typically employ SILU for this stage.
  • We then switch the activation to that used for inference for the rest of the training.
[…]
Figure 2: Swi+FT: Training loss. Most of the training is carried out with SILU, with $\alpha$=5%, 10% and 20% of the final steps using RELU. Note the loss spike when we switch the activation. The model rapidly recovers and converges to a regime where RELU is performing well while providing sparsity. This strategy needs to be combined with StochA to provide good models operating with RELU.
3.2 Stochastic activation: StochA
A stochastic function, parametrized by a random variable $\omega$, is a function
$y=\Psi(x,\omega)$ (2)
that maps inputs $x\in\mathbb{R}$ to output $y\in\mathbb{R}$ with randomness involved. The dependence on $\omega$ emphasizes that the outcome depends on an underlying probability space. In that sense, the function $\Psi(\cdot,\omega)$ is deterministic for each realization of $\omega$, but is stochastic overall. In particular, we consider the case depicted in Figure 1, where $\omega\sim\mathrm{Bernoulli}(p)$ is a binary random variable parametrized by a parameter $p$: $\omega\in\{0,1\}$ such that $\mathbb{P}(\omega=1)=p$ and $\mathbb{P}(\omega=0)=1-p$. In that case, the stochastic function $\Psi_{p}(\cdot)$ is defined such that
$\textrm{if }x<0,\ \Psi_{p}(x)=(1-\omega)\times\texttt{RELU}(x)+\omega\times\texttt{SILU}(x),$ (3)
which amounts to randomly selecting between the RELU and SILU activations for $x<0$. If $x\geq 0$ we choose either $\Psi_{p}(x)=x$ or $\Psi_{p}(x)=\texttt{SILU}(x)$, see the baseline paragraph below.
This strategy ensures that the network is compatible with two regimes. The first one, drawn with probability 1p1-p, is the inference-time mode, where we prepare the network to employ RELU during generation, in order to exhibit sparsity. The second mode aims to facilitate optimization during training. The choice of the SILU activation is motivated by the regular deterministic gated design by (Shazeer, 2020) adopted by most state-of-the-art LLMs.
Notation
To specify an activation, we separately define the function for the positive and negative range of inputs. For example R-S+ means that RELU is used for the negative range and SILU for the positive. When StochA is used, we indicate [S|R]-S+, which means that for the negative range, we sample SILU with probability $p$ and RELU with probability 1p1-p.
Baselines
The two natural baselines are the deterministic functions SILU and RELU. We also introduce two non-stochastic baselines in order to disentangle the effect that could come from combining SILU and RELU separately in the positive and negative domain: these baselines are denoted by S-R+ and S-R+.
$x<0$ $x\geq 0$
RELU 0 $x$
SILU $x\cdot\sigma(x)$ $x\cdot\sigma(x)$
R-S+ 0 $x\cdot\sigma(x)$
R+S- $\cdot\sigma(x) xx$
Discussion
The stochastic strategy resembles activation dropout (Srivastava et al., 2014), which can be regarded as a particular case of our method where one of the activations is the null function. However, the objective of dropout is to avoid overfitting. Our motivation is closer to Quantization-aware training (Jacob et al., 2017), more specifically, to the QuantNoise strategy of Fan et al. (2020), where the model is pretrained to observe both the train-time (unquantized) and the inference-time (quantized) models. In QuantNoise, using these two modes during training time ensure both the proper optimization, without suffering the flat gradient inherent to quantization, while reducing the discrepancy that results from the late-quantization of the model weights.
Alternative construction of a stochastic activation.
An alternative construction is to randomly select between the identity function $x\mapsto x$ and the constant zero function $x\mapsto 0$ with a sigmoidal probability $\sigma(x)$. As a result, in expectation this function is given by
$\mathbb{E}[\mathrm{sa(x)}]=(1-\sigma(x))\cdot 0+\sigma(x)\cdot x=\sigma(x)\cdot x,$ (4)
where we recognize the $\texttt{SILU}(x)$ function. While the simplicity of this construction is mathematically appealing, our preliminary experiments revealed that it does not work very well.
3.3 Inference-time strategies and evaluation
At test time, we evaluate and analyze models trained with Swi+FT and/or StochA as follows:
RELU at test time.
This is how we can enable sparsity. The corresponding evaluations therefore measure the performance on benchmarks when using this activation at test time.
Exploiting sparsity
On an input $x\in\mathbb{R}^{D}$, the gated FFN computes:
$y=W_{2}\times(\texttt{RELU}(W_{1}\times x)\odot(W_{3}\times x))\textrm{ with }W_{1},W_{3}\in\mathbb{R}^{N\times D}\textrm{ and }W_{2}\in\mathbb{R}^{D\times N},$ (5)
assuming column vectors and noting $\odot$ the element-wise multiplication. When the activation $\texttt{RELU}(W_{1}\times x)$ has a fraction $s$ of zeros, the multiplications by $W_{2}$ and $W_{3}$ can exploit this sparsity: the baseline of $3ND$ FLOPS reduces to (32s)ND(3-2s)ND.
Note that exploiting the sparsity to increase the computational throughput is not straightforward. At training time, the runtime is dominated by matrix-matrix multiplications, where even a 90% sparsity rate is not guaranteed to yield efficiency gains. At inference time with one prompt at a time, the bottleneck is the memory access used during matrix-vector multiplications. When $W_{2}$ is stored by rows and $W_{3}$ by columns, the sparsity can be exploited to avoid a fraction $s$ of the memory reads, that are contiguous. This implementation nearly yields the expected speedup (see experimental section).
Stochastic activation at test time.
The following only applies to the StochA strategy: we evaluate the performance when if leverage the randomness at test time, i.e., in this case, we do not use RELU. This choice has two interests:
  1. 1.
    To quantify the effect of the activation discrepancy between train and test.
  2. 2.
    To generate multiple outputs from the same prompt with the randomness of StochA.
For the second usage, the standard way to generate multiple outputs from the same prompt is to replace the greedy decoding with a random sampling of the token from its probability distribution. This sampling can be tuned by setting a softmax temperature $T$ which adjusts between completely uniform sampling ($T\rightarrow\infty$) and strict maximum sampling ($T\rightarrow 0$). In both cases, we keep the one generated output with the highest normalized log likelihoods, i.e., the per-token average log-likelihood, as predicted by the model.
4 Experiments with large language models
4.1 Experimental setting
Model architecture
We train dense decoder-only models. The transformer blocks use grouped-query attention (Ainslie et al., 2023). These models use RMSNorm (Zhang & Sennrich, 2019) with prenormalization, rotary positional positional encoding (RoPE) (Su et al., 2021) with $\theta=500000$ and train with document causal masking. We use the SILU activation (Shazeer, 2020) for the SILU baseline. The structure of our LM1.5B and LM3B models is detailed in Table 4 in Appendix A.
Training hyper-parameters
We train the models with AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}$ = $0.9$, $\beta_{2}$ = $0.95$, learning rate of $lr$ = $3\times 10^{-3}$, weight decay of $0.1$, and gradient clipping at norm $1.0$. After 2000 steps of linear warm-up, we use a cosine decay learning rate schedule with peak learning rate $8\times 10^{-4}$ and decay by a factor of $1/100$ over the training horizon.
Tokenizer
We use the Llama3 (Dubey et al., 2024) tokenizer, which is a fast Byte-Pair Encoding tokenizer implemented with TikToken.2 The vocabulary contains 128 000 regular tokens as well as 256 reserved tokens.
Pre-training
We pre-train the LM1.5B and LM3B models with 47B and 80B tokens, respectively, from a diverse collection of mostly English natural language and coding data. We use a batch size of 1M tokens and a context length of 8192 tokens.
Evaluation Benchmarks
We employ two types of benchmarks for zero or few-shot evaluation, which we describe in more detail in Appendix C and Table 6. The first type is code generation tasks: HumanEval+ (Liu et al., 2023) and MBPP (Chen et al., 2021). The second type consists of common sense and general reasoning: HellaSWAG(Zellers et al., 2019), ARC(Clark et al., 2018), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2020), NQ (Kwiatkowski et al., 2019), RACE (Lai et al., 2017), TQA (Joshi et al., 2017) and GSM8K (Cobbe et al., 2021).
4.2 Performance analysis of Swi+FT and StochA with RELU at inference time
Training activation LM1.5B LM3B
$x<0$ $x>0$ $p$ train val val train val val
Activation pp 1p1-p Swi+FT RELU StochA RELU StochA
SILU SILU SILU - 2.105 2.122 1.966 1.974
RELU RELU RELU - 2.140 2.161 2.027 2.043
S-R+ SILU RELU - 2.101 2.124 1.970 1.980
R-S+ RELU SILU - 2.123 2.151 2.016 2.033
[S|R]-R+ SILU RELU RELU 0.3 2.120 2.363 2.146 1.993 2.257 2.006
[S|R]-R+ SILU RELU RELU 0.5 2.120 2.507 2.145 1.990 2.889 1.999
[S|R]-S+ SILU RELU SILU 0.3 2.115 2.305 2.143 1.987 2.257 1.996
[S|R]-S+ SILU RELU SILU 0.5 2.115 2.530 2.143 1.984 2.753 1.995
[S|R]-R+ SILU RELU RELU 0.3 2.123 2.141 2.251 1.988 1.998 2.177
[S|R]-R+ SILU RELU RELU 0.5 2.129 2.148 2.307 1.989 2.002 2.306
[S|R]-S+ SILU RELU SILU 0.3 2.120 2.138 2.221 1.982 1.992 2.103
[S|R]-S+ SILU RELU SILU 0.5 2.125 2.144 2.301 1.985 1.994 2.234
Table 1: Losses: train is computed over the last 500 steps of the training loss of LM1.5B, val is measured after training on a different set set of text and code using the RELU activation or StochA, i.e., the same activation used at train time (possibly deterministic). If Swi+FT is enabled, we switch to RELU for the last $5\%$ steps. : for the deterministic baselines SILU, S-R+ and S+R-, we do the inference with the same activation used at train-time (not RELU).
In this section we analyze the effect of our proposal when using RELU at test time. In Appendix B, we provide a complementary analysis of the sparsity. Depending on the setting, the average rate of 0s can be higher than 90%, when using the RELU at test time.
Cross-entropy performance.
Table 1 provides the impact on the training and validation losses of multiple choices with the LM1.5B and LM3B models. We observe that the training loss using stochastic activation at train time is lower than that obtained with RELU. However the validation entropy is not competitive per se, due to the remaining train-inference discrepancy of activation. This is solved by Swi+FT: switching to the RELU activation function and fine-tuning for the last 5% or 10% steps percentage of the training steps drastically boosts the test-time inference. These results outperform the results obtained with regular RELU training, while using the same activation at test time.
Fast inference with RELU sparsity
The activation sparsity can be exploited to avoid fetching 90% of the matrices $W_{1}$ and $W_{3}$ of the FFN (see Eq. 5 ands detailled sparsity rate in Appendix B). This has a direct benefit on CPU, see Figure 3: the 90% sparsity provides a 65% speedup. On GPU the additional challenge is to make the computation sufficiently predictable to balance the load between CUDA threads.
Figure 3: Total inference time for 1 token on CPU, as a function of the activation sparsity, with a LM3B model trained with Swi+FT. The “other operations" include the attention layers (they are not dominant because the generation is limited to 200 tokens), the normalization and the execution overheads. At 90% sparsity the speedup is $\times$1.65. The timings are measured on a single core of a Xeon 8462Y+ machine.
Complementary of StochA and Swi+FT with RELU finetuning
Figure 4 shows the training loss when using StochA and Swi+FT jointly. When switching the activation to RELU, we observe a spike in the loss, but the optimization rapidly recovers, and then converges to a model having a better performance than the one produced when training with RELU from scratch. In contrast, in the case where we employ Swi+FT alone (Figure 2), fine-tuning with RELU after pretraining with SILU is not enough to obtain performance improvements. This shows that both approaches are complementary.
final loss
$\alpha$ train val (RELU)
0 2.115 2.530
0.05 2.125 2.144
0.10 2.119 2.145
0.20 2.126 2.155
Figure 4: Training loss with Swi+FT and StochA: [S|R]-S+ activation with $p=0.5$ for $\alpha$$5\%$, $10\%$ and $20\%$, relative to RELU and SILU. Note that this plot this is zoomed in relative to Figure 2.
The impact of $\alpha$
can also be seen in Figure 4, where we compare the final loss reported for different fine-tuning values. In that case, setting $\alpha=0.1$ is a good trade-off since the corresponding loss is lower than the RELU loss.
4.3 Performance on downstream tasks
Detailed results per benchmark
Table 2 reports the results for the standard code generation, common sense and general reasoning benchmarks detailed in Appendix C. We consider multiple StochA models using few-shot or zero shot prompting, see Table 6 in the appendix for more details for each case. Observe first that the model with SILU is significantly better that a regular model with RELU. However, our models trained with StochA are slightly better or on par with SILU: either the model fine-tuned with Swi+FT and using RELU at inference time, or even the model that uses StochA at test time.
$\downarrow$ Benchmark/metric LM1.5B LM3B
baselines (a) Swi+FT (b) StochA baselines (a) Swi+FT (b) StochA
train activation $\rightarrow$ SILU RELU [S|R]-S+ [S|R]-S+ SILU RELU [S|R]-S+ [S|R]-S+
inference activation $\rightarrow$ SILU RELU RELU [S|R]-S+ SILU RELU RELU [S|R]-S+
hellaswag/acc_char 0.585 0.561 0.574 0.576 0.684 0.633 0.671 0.678
winogrande/acc_char 0.593 0.571 0.568 0.568 0.657 0.615 0.630 0.620
arc_easy/acc_char 0.568 0.562 0.600 0.562 0.675 0.642 0.679 0.671
arc_challenge/acc_char 0.313 0.286 0.331 0.314 0.390 0.348 0.396 0.376
piqa/acc_char 0.732 0.720 0.724 0.720 0.767 0.751 0.765 0.761
obqa/acc_char 0.346 0.340 0.378 0.340 0.390 0.380 0.384 0.408
race.middle/acc_char 0.518 0.516 0.509 0.498 0.565 0.538 0.559 0.549
race.high/acc_char 0.382 0.379 0.372 0.375 0.414 0.402 0.407 0.416
human_eval_plus/pass@1 0.073 0.067 0.049 0.055 0.128 0.110 0.128 0.116
mbpp/compiles@1 0.978 0.970 0.960 0.980 0.992 0.980 0.990 0.982
tqa/f1 0.243 0.217 0.229 0.232 0.351 0.293 0.327 0.342
nq/f1 0.123 0.107 0.121 0.113 0.169 0.146 0.170 0.145
average performance 0.454 0.441 0.451 0.444 0.515 0.486 0.509 0.505
Table 2: Performance per benchmark of the RELU and SILU (a) baselines for LM1.5B and LM3B compared to (b) models with StochA and Swi+FT at train time and RELU at test time, and (c) models with StochA at train and test time. We use the model with the best perplexity on val namely $p=0.3,\alpha=0.05$ for Swi+FT and $p=0.5$ for StochA.
Figure 5: Swi+FT: analysis of the fine-tuning rate $\alpha$. Average performance over the benchmarks as a function of the percentage $\alpha$ of steps for which we switch to the RELU activation at the end of training. We use RELU at inference time.
Performance when varying $\boldsymbol{\alpha}$ with Swi+FT.
Figure 5 showcases that we can slightly surpass the SILU baseline if we first use a stochastic activation function during the LM1.5B model training and then switch to the RELU activation function for the last $\alpha\%$ of the training steps, for $\alpha\in\{5\%,10\%,20\%\}$. The best performance is typically obtained with $\alpha=5\%$ or $\alpha=10\%$ for the LM1.5B model.
4.4 Exploiting StochA at test time
Effectiveness of StochA a test time.
In Table 1, in addition to the results with RELU at test time, we also report the train and validation losses obtained when employing StochA at test time. We observe that (1) using stochastic activations for inference works surprisingly well in spite of the randomness. The results are between RELU and SILU in most configurations; (2) When using StochA at test time, there is no need to fine-tune the model with Swi+FT. This is expected since this strategy is intended to decrease the discrepancy with the test-time activation choice. Table 3 shows that the average benchmark performance generally increases when the stochastic mix approaches SILU. Therefore, StochA is primarily useful as a way to generate multiple outputs for the same prompt.
$p$ LM1.5B LM3B
0.1 (R-S+) 0.215 0.222
0.3 0.443 0.504
0.5 0.426 0.505
0.7 0.453 0.495
1.0 (SILU) 0.454 0.515
Table 3: StochA: Impact on benchmarks performance (avg) as a function of the StochA $p$ for [S|R]-S+ used at test time. The case $p=0$ corresponds to R-S+ while $p=1$ corresponds to the baseline SILU. The performance increases with more SILU in the mix. However, the stochasticity can be used to increase the generation diversity.
NQ TQA
Refer to caption Refer to caption
Figure 6: Comparison of ways to generate diverse responses in Q&A benchmarks: by varying the softmax temperature or by varying the [S|R]-S+ pp. The multiple generations are scored by normalized log likelihoods and the best generation is evaluated with an F1 score (y-axis).
Diversity of generations ablations.
Figure 6 shows how stochastic generation compares to temperature sampling to generate diverse outputs, see Section D for examples. We plot these on TQA and NQ, whose results are least noisy. For the best settings, the curves are increasing, which means that (1) the generations are indeed diverse and (2) that the normalized log likelihoods are a suitable scoring function. The StochA activations yield a higher performance than vanilla temperature sampling on NQ but are sub-par for TQA.
5 Conclusion
This paper has introduced a novel stochastic activation that preserves the performance of a non-sparse activation, such as SILU, while better adjusting to the behavior of a sparse one, such as RELU, at test time. This improves the inference times for the FFN layers of a transformer, translating into a speedup of typically $\times 1.65$ for the FFN processing on CPUs while almost preserving the accuracy of the non-sparse SILU activation. Finally, we explore how stochastic activations can be leveraged at test time to improve diversity in model generations.
Appendix
In this appendix, we add some details that did not fit in the main paper. Appendix A gives some details about the models used in the paper. Appendix B gives additional details about the sparsity of the activations produced by the models that we use. Appendix C gives the references for all the benchmarks used in the experiments. The final Appendix D gives some examples of multiple answers generated by StochA.
Appendix A Architecture detail
The model architectures we use in the paper are loosely inspired by Llama 3. Table 4 summarizes their main parameters.
Parameter LM1.5B LM3B
Number of parameters 1.5B 3B
Layers 28 36
Hidden dimension 1536 2048
Intermediate dimension 8960 11008
Number of attention heads 12 16
Number of key-value heads 2 2
Table 4: LM1.5B and LM3B model parameters
Appendix B Sparsity Analysis
train inference Swi+FT StochA: p Swi+FT: $\alpha$ sparsity (%)
baselines SILU SILU 00000.0002
RELU RELU 94.8
Swi+FT SILU RELU 0.05 79.9
StochA [S|R]-S+ RELU 0.3 0.05 88.5
+Swi+FT [S|R]-S+ RELU 0.5 0.05 86.5
[S|R]-S+ RELU 0.7 0.05 84.6
Table 5: We report the rate of the 0-valued activation after the SILU (not sparse) and RELU activation, when training with the LM1.5B model with the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$.
Pre-activation inputs distribution dead neurons in the $W_{1}$ layer
Refer to caption Refer to caption
Figure 7: Left: empirical cumulative distribution functions for the inputs to the activation functions when training the LM1.5B model with one of the following activations: SILU, RELU, Swi+FT fine-tuning and the combination of StochA with Swi+FT for different values of $p$. Right: fraction of near-zero rows of the $W_{1}$ matrix, indicating useless neurons for two types of activations.
In Table 5, we report the sparsity rates resulting from the gating in the FFN layer, when using the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$. In all cases, except for the SILU baseline, we use RELU at test time. We observe that Swi+FT fine-tuning with RELU brings back a lot of sparsity (80%), yet the model is not competitive. The combination of StochA with Swi+FT leads to models offering a high sparsity degree by using RELU at test time and simultaneously achieving competitive performance, see Tables 1 and 2.
In Figure 7 (left), we observe that $P(X\leq 0)$ is the highest for RELU since most of its mass is concentrated around zero as in Table 5. For $x<0$, the probability $P(X\leq x)$ is lower for SILU than for both Swi+FT and when we combine StochA with Swi+FT for different values of $p$. This shows that combining StochA with Swi+FT $P(X\leq x)$ enables us to control how much mass we assign to negative values of $x$ depending on $p$.
The sudden jump in the CDF for RELU indicates that there are many activations that are exactly 0 on input of RELU. When investigating the rows of linear layer $W_{1}$ (Eq 5) in Figure 7 (right), it turns out a large fraction of these are near 0 (i.e. their norm is less than 1/1000 the average row norm of the matrix). This means that these rows are unused: they are “dead neurons”. Fortunately, with our Swi+FT approach, that uses SILU at pre-training, these dead neurons disappear. This explains why our approach can obtain significantly better results with RELU than the vanilla training of the model with the RELU activation employed from scratch.
Appendix C Benchmarks
Code generation
We use two benchmarks that evaluate the code generation capabilities of AI models: HumanEval+ and MBPP.
  • The HumanEval+ (Liu et al., 2023) benchmark is an extension of HumanEval (Chen et al., 2021), which is designed to evaluate the functional correctness of code generated by AI models.
  • MBPP (Austin et al., 2021) is designed to evaluate the code generation abilities of AI models, particularly for Python programming tasks.
Common sense and general reasoning
We use benchmarks consisting of question-answer or multiple-choice questions designed to evaluate the commonsense reasoning abilities of AI models, particularly in the context of natural language understanding: HellaSWAG, ARC, PIQA, OBQA, Winogrande, NaturalQuestions, RACE, TQA and GSM8K.
  • HellaSWAG (Zellers et al., 2019) consists of multiple-choice questions where each question contains a short context (a sentence or paragraph) followed by four possible continuations. Only one continuation is correct and makes sense given the context.
  • The AI Reasoning Challenge (ARC) (Clark et al., 2018) benchmark consists of multiple-choice science questions typically found in elementary and middle school exams. The ARC questions require a mix of factual knowledge, commonsense reasoning, and multi-step inference.
  • The Physical Interaction Question-Answering (PIQA) (Bisk et al., 2020) benchmark consists of multiple-choice questions about how to accomplish simple physical tasks. Each question presents a short scenario and two possible solutions; only one is physically plausible.
  • The OpenBook QA (OBQA) (Mihaylov et al., 2018) benchmark consists of 6,000 multiple-choice questions based on elementary science facts. Each question is designed to require combining a provided “open book” science fact with additional commonsense or general knowledge.
  • WinoGrande (Sakaguchi et al., 2020) benchmark consists of multiple-choice questions. Each question presents a sentence with a pronoun and two possible antecedents; the task is to choose the correct referent for the pronoun.
  • Natural Questions (NQ) (Kwiatkowski et al., 2019) benchmark is designed to evaluate the ability of an AI model to answer real user questions using information from Wikipedia.
  • The Reading Comprehension from Examinations (RACE) (Lai et al., 2017) benchmark consists of passages and multiple-choice questions to assess how well AI models can comprehend and reason about written passages.
  • The Trivia QA benchmark (TQA) (Joshi et al., 2017) is a reading comprehension dataset that pairs trivia questions with evidence documents from which answers can be derived.
  • The Grade School Math 8k (GSM8k) (Cobbe et al., 2021) benchmark is a dataset of 8,500 high-quality, linguistically diverse grade school math word problems. The benchmark is designed to test multi-step mathematical reasoning capabilities in language models.
Benchmark Metric Few Shot Type
hellaswag acc_char 0 choice
winogrande acc_char 0 choice
arc_easy acc_char 0 choice
arc_challenge acc_char 0 choice
piqa acc_char 0 choice
obqa acc_char 0 choice
race.middle acc_char 0 choice
race.high acc_char 0 choice
human_eval_plus pass@1 0 generation
mbpp compiles@1 3 generation
tqa f1 5 generation
nq f1 5 generation
Table 6: Summary of benchmarks used for evaluation. We few-shot prompt the LM1.5B and LM3B models. The metric depends on the benchmark as well as the average performance.
Appendix D Examples of various sequences generated with StochA
Using a pre-trained LM3B model with StochA, we can use the StochA activation at test time to generate multiple predictions by leveraging the randomness from the activation function. We provide two examples of the generations obtained using questions from the TQA benchmark (Joshi et al., 2017):
question and generated number of NLL F1
accepted answers answers times score score
When did ibuprofen become available over the counter? A: 1983 A: 1984 1988 4 1.002 0
1982 11 1.004 0
1980 1 1.030 0
1995 2 1.019 0
2001 1 1.017 0
1983 1 1.024 1
Who played michael jackson in jackson 5 movie? A:alex burrall A:abolade david olatunde A:wylie draper A:jason weaver Michael Jackson 19 0.948 0
Michael jackson 1 0.792 0
Where is the meridian that is opposite the prime meridian located? A:180th meridian A:antimeridian Greenwich 9 0.896 0
Greenwich, England 11 0.551 0
The cold dry winds that blow over northern india in winter are called? A:northeast monsoon A:retreating monsoon A:northeast monsoon or retreating monsoon Siberian 7 1.099 0
Siberian winds 6 0.970 0
Ganges 1 1.617 0
monsoon 2 1.051 2/3
katabatic winds 3 0.738 0
Ganga 1 1.643 0
Where can you find dna in the body? A:chromosomes in cell A:inside cell nucleus Blood 6 1.262 0
Hair 12 1.393 0
Mitochondria 1 0.700 0
In the nucleus 1 1.437 2/5
Who is often associated with printing the first book using moveable type in germany? A:johannes gutenberg Johannes Gutenberg 20 0.205 1
Who won the womens 2017 ncaa basketball tournament? A:south carolina Kentucky 15 1.852 0
Kentucky Wildcats 1 1.304 0
North Carolina 4 1.342 1/2
Country with most olympic gold medals all time? A:united states USA 14 0.818 0
United States 6 0.614 1
The atomic number of indium which belongs to 5th period is? A:49 5 14 0.579 0
49 5 0.650 1
84 1 0.695 0
Who appoints the members of the board of governors of the federal reserve? A:president The president 16 0.602 1
The president of the United States 4 0.431 2/5
What age do you need to be to buy a bb gun? A:18 14 17 0.660 0
10 2 0.680 0
18 1 0.728 1
What genre is the magic tree house books? A:childrens historical fantasy Fantasy 14 1.038 1/2
Children’s fiction 1 1.051 2/5
Children’s 3 1.245 1/2
Children’s books 2 1.067 2/5
What is the name of the skin between your nostrils? A:nasal septum A:septum Nasal septum 17 0.707 1
Nasion 3 1.224 0
Table 7: Example generations. For each question, we indicate the ground-truth answers (from the dataset). We generate 20 answers per question with [S|R]-S+, pp=0.7. We list the de-duplicated answers, with the NLL score (used to sort the results) and the F1 score (used to evaluate the result, it is computed as the intersection of bags-of-words).

### v2

@@
 As the demand for more efficient and powerful models grows, researchers are increasingly focused on optimizing their training processes to balance performance with resource constraints.
-The majority of the LLM parameters are in the Feed-Forward Network (FFN) layers, where they memorize the training data.
+The majority of the LLM parameters are in the Feed-Forward Network (FFN) layers, where they primarily serve to store and recall information from the training data.
 FFNs are two linear layers separated by an activation function, and sometimes an additional linear layer that serves as a gating operation.
-The activation is a $\mathbb{R}\rightarrow\mathbb{R}$ non-linear function.
+The activation is a non-linear function from $\mathbb{R}$ to $\mathbb{R}$.
 In this context, the choice of activation function plays a crucial role for both the model’s expressivity and efficiency.
-The simplest activation is RELU (Rectified Linear Unit), that allows positive values to pass through and forces negative inputs to zero.
+The simplest choice of activation is RELU (Rectified Linear Unit), that allows positive values to pass through and forces negative inputs to zero.
 RELU is sparsity-inducing, since, on average, half of its outputs are zero (in practice significantly more).
-Within a two-layer Multilayer Perceptron, this means that inference on the second layer is a matrix-sparse vector multiplication, so it can be implemented with fewer FLOPs than a matrix-dense vector multiplication.
+Within a two-layer Multilayer Perceptron, this means that inference on the second layer is a matrix-sparse vector multiplication, hence, it can be implemented with fewer FLOPs than a matrix-dense vector multiplication.
 Note that effectively improving the runtime with this sparsity pattern remains challenging.
@@
 In this work, we consider two ways to approach this problem.
-The first approach is activation fine-tuning, denoted Swi+FT: we pre-train the model with an activation that facilitates efficient large language model optimization, then we change the activation to RELU and adapt the model by fine-tuning it further.
+The first approach is activation fine-tuning, denoted by Swi+FT: we pre-train the model with an activation that facilitates efficient large language model optimization, then, we change the activation to RELU and adapt the model by fine-tuning it further.
 Our second approach, referred to as StochA (stochastic activations), is a novel technique that randomly selects between multiple activations, either at train or test time.
@@
 In summary, this paper makes the following contributions:
-  • We introduce and analyze two strategies that employ different activation functions at training and inference time, namely Swi+FT and StochA.
+  • We introduce two strategies that employ different activation functions at training and inference time, namely Swi+FT and StochA.
 Both are complementary and make it possible to use activations at inference time that differ from those employed during pre-training.
   • We produce RELU-based models that are much better than those obtained with regular training, i.e., our methods significantly outperform training with RELU only.
-  • We show that stochastic activations, when used at inference time, provides an alternative way to generate diverse sequences compared to traditional temperature sampling or other variants.
+  • We show that stochastic activations, when used at inference time, provide an alternative way to generate diverse sequences compared to traditional temperature sampling or other variants.
 2 Related work
 Standard activation functions
-Activations are at the core of deep learning, in that they are needed to depart from defining a linear function with limited expressivity.
-While early neural network architectures were inspired by logistic regression, such as sigmoidal and tanh activations, many activation functions have been evaluated for the Feed Forward layers (FFNs) of transformers.
+Activations are at the core of deep learning, they are fundamental to enable deep learning models to go beyond linear function with limited expressivity.
+While early neural network architectures were inspired by logistic regression, such as sigmoidal and tanh activations, many activation functions have been evaluated for the FFN layers of transformers.
 Vaswani et al.
-(2017) used RELU (Glorot et al., 2011).
+(2017) used RELU (Glorot et al., 2011) initially.
 However, using the RELU activation function leads to some neurons getting stuck in the negative region.
-As a consequence, they stop learning entirely, since the gradient is zero for negative inputs, and their weights do not get updated.
+As a consequence, models stop learning entirely, since the gradient is zero for negative inputs, and their weights do not get updated.
 In contrast, Touvron et al.
@@
 Lee et al.
-(2022) propose an activation, called Adaptive Swish (ASH), that uses stochastic sampling of the top-k percentile elements.
-It is a generalization of the Swish (Ramachandran et al., 2018) activation function, that uses adaptive thresholding and selects only the values in the top percentiles and sets these to zero otherwise.
-The threshold is not fixed but depends on the distribution of the input making it an example of a stochastic activation.
+(2022) propose the Adaptive Swish(ASH) activation, which uses stochastic sampling of the top-k percentile elements.
+It generalizes the Swish (Ramachandran et al., 2018), which uses adaptive thresholding to select the values in the top percentiles and is set to zero otherwise.
+It is an example of a stochastic activation.
 Dropout and structured Dropout variants
@@
 LayerDrop (Fan et al., 2019) randomly drops entire layers during training, hence, it encourages the model to be robust to missing layers.
-At inference time, some layers can be pruned, trading off between speed and accuracy as needed.
+At inference time, some layers can be pruned, trading off speed and accuracy as needed.
 While the method does not make the model sparse in the usual sense, it induces structured sparsity in the computation graph during training.
@@
 Figure 2: Swi+FT: Training loss.
-Most of the training is carried out with SILU, with $\alpha$=5%, 10% and 20% of the final steps using RELU.
+Most of the training is carried out with SILU, with $\alpha*100\%$=5%, 10% and 20% of the final steps using RELU.
 Note the loss spike when we switch the activation.
 The model rapidly recovers and converges to a regime where RELU is performing well while providing sparsity.
-This strategy needs to be combined with StochA to provide good models operating with RELU.
 3.2 Stochastic activation: StochA
@@
 $\textrm{if }x<0,\ \Psi_{p}(x)=(1-\omega)\times\texttt{RELU}(x)+\omega\times\texttt{SILU}(x),$ (3)
-which amounts to randomly selecting between the RELU and SILU activations for $x<0$. If $x\geq 0$ we choose either $\Psi_{p}(x)=x$ or $\Psi_{p}(x)=\texttt{SILU}(x)$, see the baseline paragraph below.
+which corresponds to randomly selecting between the RELU and SILU activations for $x<0$. If $x\geq 0$ we choose $\Psi_{p}(x)=x$ or $\Psi_{p}(x)=\texttt{SILU}(x)$, see the baselines description below.
 This strategy ensures that the network is compatible with two regimes.
@@
 R-S+ 0 $x\cdot\sigma(x)$
-R+S- $\cdot\sigma(x) xx$
+S-R+ $\cdot\sigma(x) xx$
 Discussion
@@
 At test time, we evaluate and analyze models trained with Swi+FT and/or StochA as follows:
-RELU at test time.
+RELU at test time
 This is how we can enable sparsity. The corresponding evaluations therefore measure the performance on benchmarks when using this activation at test time.
@@
 This implementation nearly yields the expected speedup (see experimental section).
-Stochastic activation at test time.
-The following only applies to the StochA strategy: we evaluate the performance when if leverage the randomness at test time, i.e., in this case, we do not use RELU. This choice has two interests:
+Stochastic activation at test time
+The following only applies to the StochA strategy: we evaluate the performance by leveraging the randomness at test time, i.e., in this case, we do not use RELU.
+This choice is interesting for two reasons:
   1. 1.
@@
 Evaluation Benchmarks
-We employ two types of benchmarks for zero or few-shot evaluation, which we describe in more detail in Appendix C and Table 6.
+We employ two types of benchmarks for zero or few-shot evaluation, which we describe in more detail in Appendix G and Table 8.
 The first type is code generation tasks: HumanEval+ (Liu et al., 2023) and MBPP (Chen et al., 2021).
@@
 [S|R]-S+ SILU RELU SILU 0.5 2.125 2.144 2.301 1.985 1.994 2.234
-Table 1: Losses: train is computed over the last 500 steps of the training loss of LM1.5B, val is measured after training on a different set set of text and code using the RELU activation or StochA, i.e., the same activation used at train time (possibly deterministic).
+Table 1: The train loss is computed over the last 500 steps of the training of LM1.5B, the val loss is measured after training, on a different set of text and code, using the RELU activation or StochA, i.e.
+the same activation used at train time (possibly deterministic).
 If Swi+FT is enabled, we switch to RELU for the last $5\%$ steps.
@@
 Depending on the setting, the average rate of 0s can be higher than 90%, when using the RELU at test time.
-Cross-entropy performance.
-Table 1 provides the impact on the training and validation losses of multiple choices with the LM1.5B and LM3B models.
+Cross-entropy performance
+Table 1 summarizes the impact on the training and validation losses of multiple choices with the LM1.5B and LM3B models.
 We observe that the training loss using stochastic activation at train time is lower than that obtained with RELU.
 However the validation entropy is not competitive per se, due to the remaining train-inference discrepancy of activation.
-This is solved by Swi+FT: switching to the RELU activation function and fine-tuning for the last 5% or 10% steps percentage of the training steps drastically boosts the test-time inference.
+This is solved by Swi+FT: switching to the RELU activation function and fine-tuning for the last 5% or 10% steps of the training steps drastically boosts the test-time inference.
 These results outperform the results obtained with regular RELU training, while using the same activation at test time.
 Fast inference with RELU sparsity
-The activation sparsity can be exploited to avoid fetching 90% of the matrices $W_{1}$ and $W_{3}$ of the FFN (see Eq.
-5 ands detailled sparsity rate in Appendix B).
-This has a direct benefit on CPU, see Figure 3: the 90% sparsity provides a 65% speedup.
-On GPU the additional challenge is to make the computation sufficiently predictable to balance the load between CUDA threads.
-Figure 3: Total inference time for 1 token on CPU, as a function of the activation sparsity, with a LM3B model trained with Swi+FT.
-The “other operations" include the attention layers (they are not dominant because the generation is limited to 200 tokens), the normalization and the execution overheads.
+[…]
+CPU (one core of Xeon 8462Y+) GPU (H100 80GB HBM3)
+Refer to caption Refer to caption
+Figure 3: Inference times for 1 token, as a function of the activation sparsity, with a LM3B model trained with Swi+FT.
+For the CPU we indicate the total inference time, including “other operations": the attention layers (these are not dominant because the generation is limited to 200 tokens), the normalization and the execution overheads.
 At 90% sparsity the speedup is $\times$1.65.
-The timings are measured on a single core of a Xeon 8462Y+ machine.
-Complementary of StochA and Swi+FT with RELU finetuning
+For the GPU, we measure only the FFN inference time.
+Note that the impact on inference is less important because at this speed and model scale the overheads are dominant.
+We also indicate the minimum runtime that could be obtained given the GPU’s memory bandwidth (roofline model).
+Absence of dead neurons with Swi+FT
+Figure 7(left) shows a sudden jump in the RELU CDF, which indicates there are many activations that are exactly zero on input of RELU.
+The RELU plot in Figure 7(right) shows that a large fraction of the rows of the linear layer $W_{1}$ (Eq 5) are near zero (i.e.
+their norm is less than 1/1000 the average row norm of the matrix).
+This means that these rows are unused: they are “dead neurons”.
+In contrast, the Swi+FT plot in Figure 7(right) we observe that using SILU at pre-training prevents these dead neurons appearing.
+This explains why our approach obtains significantly better results with RELU than the vanilla training of the model using the RELU from scratch.
+Complementarity between StochA and Swi+FT when fine-tuned with RELU
 Figure 4 shows the training loss when using StochA and Swi+FT jointly.
-When switching the activation to RELU, we observe a spike in the loss, but the optimization rapidly recovers, and then converges to a model having a better performance than the one produced when training with RELU from scratch.
+When switching the activation to RELU, we observe a spike in the loss, but the optimization rapidly recovers, and converges to a model that has better performance than the one trained with RELU from scratch.
 In contrast, in the case where we employ Swi+FT alone (Figure 2), fine-tuning with RELU after pretraining with SILU is not enough to obtain performance improvements.
-This shows that both approaches are complementary.
+As a consequence, the StochA and Swi+FT approaches are complementary.
 final loss
@@
 0.20 2.126 2.155
-Figure 4: Training loss with Swi+FT and StochA: [S|R]-S+ activation with $p=0.5$ for $\alpha$$5\%$, $10\%$ and $20\%$, relative to RELU and SILU.
-Note that this plot this is zoomed in relative to Figure 2.
-The impact of $\alpha$
-can also be seen in Figure 4, where we compare the final loss reported for different fine-tuning values.
-In that case, setting $\alpha=0.1$ is a good trade-off since the corresponding loss is lower than the RELU loss.
+Figure 4: Training loss with Swi+FT and StochA: [S|R]-S+ activation with $p=0.5$ for $\alpha*100\%$$5\%$, $10\%$ and $20\%$, relative to RELU and SILU.
+This shows that the Swi+FT strategy needs to be combined with StochA to provide good models operating with RELU compared to finetuning SILU with RELU alone (Figure 2).
+This plot this is zoomed in relative to Figure 2.
+Impact of fine-tuning with RELU for the last $\alpha*100\%$ steps
+In Figure 4, we compare the final loss reported for different values of $\alpha$. In that case, setting $\alpha=0.1$ is a good trade-off since the corresponding loss is lower than the RELU loss.
+Continuous pre-training with StochA and Swi+FT
+Starting from a pre-trained LM1.5B model with SILU, we can leverage either StochA or Swi+FT methods for continuous pretraining (CPT), see Appendix E for experimental details.
+In such case, the best strategy is to use Swi+FT, i.e., simply fine-tune with RELU the model initially trained with SILU to obtain the best trade-off between sparsity and loss (Table 6).
+This pattern is the opposite of what we observe for LMs trained from scratch, where the StochA and StochA +Swi+FT offer the best trade-off between sparsity and loss (Table 1).
+Generalizability of StochA and Swi+FT for any pair of activations
+In Appendix F, we explain how our methods can be generalized for any pair of activation functions and include an illustrative example with the (TANH,RELU) activations.
 4.3 Performance on downstream tasks
 Detailed results per benchmark
-Table 2 reports the results for the standard code generation, common sense and general reasoning benchmarks detailed in Appendix C.
-We consider multiple StochA models using few-shot or zero shot prompting, see Table 6 in the appendix for more details for each case.
-Observe first that the model with SILU is significantly better that a regular model with RELU.
+Table 2 reports the results for standard code generation, common sense and general reasoning benchmarks detailed in Appendix G.
+We consider multiple StochA models using few-shot or zero shot prompting, see Table 8 in Appendix G for more details.
+The model with SILU (topline) is significantly better that a regular model with RELU.
 However, our models trained with StochA are slightly better or on par with SILU: either the model fine-tuned with Swi+FT and using RELU at inference time, or even the model that uses StochA at test time.
 $\downarrow$ Benchmark/metric LM1.5B LM3B
-baselines (a) Swi+FT (b) StochA baselines (a) Swi+FT (b) StochA
+(a) baselines (b) Swi+FT (c) StochA (a) baselines (b) Swi+FT (c) StochA
 train activation $\rightarrow$ SILU RELU [S|R]-S+ [S|R]-S+ SILU RELU [S|R]-S+ [S|R]-S+
@@
 average performance 0.454 0.441 0.451 0.444 0.515 0.486 0.509 0.505
-Table 2: Performance per benchmark of the RELU and SILU (a) baselines for LM1.5B and LM3B compared to (b) models with StochA and Swi+FT at train time and RELU at test time, and (c) models with StochA at train and test time.
+Table 2: Performance per benchmark of (a) RELU and SILU baselines for LM1.5B and LM3B compared to (b) models with StochA +Swi+FT at train time and RELU at test time, and (c) models with StochA at train and test time.
 We use the model with the best perplexity on val namely $p=0.3,\alpha=0.05$ for Swi+FT and $p=0.5$ for StochA.
+The average performance for the topline (SILU) is in bold, second best (StochA +Swi+FT) is underlined.
 Figure 5: Swi+FT: analysis of the fine-tuning rate $\alpha$.
@@
 We use RELU at inference time.
-Performance when varying $\boldsymbol{\alpha}$ with Swi+FT.
-Figure 5 showcases that we can slightly surpass the SILU baseline if we first use a stochastic activation function during the LM1.5B model training and then switch to the RELU activation function for the last $\alpha\%$ of the training steps, for $\alpha\in\{5\%,10\%,20\%\}$.
-The best performance is typically obtained with $\alpha=5\%$ or $\alpha=10\%$ for the LM1.5B model.
+Performance when varying $\boldsymbol{\alpha}$ with Swi+FT
+Figure 5 shows that we can slightly surpass the SILU baseline average performance if we first use a stochastic activation function during the LM1.5B model training and then switch to the RELU activation for the last $\alpha*100\%$ of the training steps, for $\alpha\in\{0.05,0.1,0.2\}$.
+The best performance is obtained with $\alpha=0.05$ or $\alpha=0.1$ for the LM1.5B model.
 4.4 Exploiting StochA at test time
-Effectiveness of StochA a test time.
+Effectiveness of StochA a test time
 In Table 1, in addition to the results with RELU at test time, we also report the train and validation losses obtained when employing StochA at test time.
@@
 However, the stochasticity can be used to increase the generation diversity.
+(a) F1 scores:
 NQ TQA
 Refer to caption Refer to caption
-Figure 6: Comparison of ways to generate diverse responses in Q&A benchmarks: by varying the softmax temperature or by varying the [S|R]-S+ pp.
-The multiple generations are scored by normalized log likelihoods and the best generation is evaluated with an F1 score (y-axis).
-Diversity of generations ablations.
-Figure 6 shows how stochastic generation compares to temperature sampling to generate diverse outputs, see Section D for examples.
-We plot these on TQA and NQ, whose results are least noisy.
-For the best settings, the curves are increasing, which means that (1) the generations are indeed diverse and (2) that the normalized log likelihoods are a suitable scoring function.
-The StochA activations yield a higher performance than vanilla temperature sampling on NQ but are sub-par for TQA.
+(b) Average type-token ratio:
+Refer to caption Refer to caption
+Figure 6: Comparison of diverse response generation methods in Q&A benchmarks.
+(a) F1 scores when varying the temperature with SILU or [S|R]-S+ $p$ with temp=0=0, the generations are scored by normalized log likelihoods.
+(b) TTR when varying the temperature in temperature sampling with SILU (solid lines) vs [S|R]-S+ for different values of $p$ (dashed lines).
+Diversity of generated sequences ablation
+Figure 6 shows that leveraging the StochA activations stochasticity (combined with standard temperature sampling) generates more diverse sequences compared to using SILU with standard temperature sampling.
+See Appendix C for the experimental setup and Appendix H for qualitative examples of such generations.
+In Figure 6 (a), the curves are increasing, which means that (1) we obtain diverse generations and (2) that the normalized log-likelihood is a suitable scoring function.
+Specifically, for the NQ task, the StochA activation with $p=0.7$ and temperature zero yields consistently higher performance than standard temperature sampling with the SILU activation (for temperatures in $\{0.1,0.2,0.5\}$).
+The StochA activation with $p=0.3$ is also competitive for the number of generations up to five.
+In contrast, the results are subpar for the TQA task: standard temperature sampling with SILU activation yields higher performances for all number of generations for temperatures in $\{0.1,0.2\}$.
+In Figure 6 (b), we observe that all the models that use StochA activation functions have consistently higher average TTR for both tasks, for each temperature in $\{0,0.1,0.2,0.3\}$ compared to standard temperature sampling with SILU.
+This finding further confirms that the text diversity of generations is much better for these models.
+Token-level diversity ablation
+In Appendix D, we highlight that using StochA at inference time is equivalent to applying Monte Carlo dropout (Gal & Ghahramani, 2016) on the negative side of the input after activation.
+We then leverage this token level diversity at inference time by doing multiple forward passes per generated token using greedy decoding.
+As observed in the sequence level ablation, more diverse tokens are generated.
+However, we did not observe that this resulted in better performance if the maximum probability or majority vote selection strategies were used.
 5 Conclusion
-This paper has introduced a novel stochastic activation that preserves the performance of a non-sparse activation, such as SILU, while better adjusting to the behavior of a sparse one, such as RELU, at test time.
+This paper introduces a novel stochastic activation that preserves the performance of a non-sparse activation, such as SILU, while better adjusting to the behavior of a sparse one, such as RELU, at test time.
 This improves the inference times for the FFN layers of a transformer, translating into a speedup of typically $\times 1.65$ for the FFN processing on CPUs while almost preserving the accuracy of the non-sparse SILU activation.
@@
 Appendix
-In this appendix, we add some details that did not fit in the main paper.
-Appendix A gives some details about the models used in the paper.
-Appendix B gives additional details about the sparsity of the activations produced by the models that we use.
-Appendix C gives the references for all the benchmarks used in the experiments.
-The final Appendix D gives some examples of multiple answers generated by StochA.
+In this appendix, we add some extra details to complement the main paper.
+Appendix A gives architectural details for the models used in the paper.
+Appendix B gives additional details of the sparsity produced by the models that use the StochA and Swi+FT activations.
+Appendix C provides the experimental details for the diversity of generations ablation.
+Appendix D provides the experimental details for the multiple forward passes per generated token ablation.
+Appendix E gives some details for the CPT experiment with StochA and Swi+FT.
+Appendix F discusses how to generalize the StochA and Swi+FT methods for any pair of non-sparse and sparse activations.
+Appendix G gives the references for all the benchmarks used in the experiments.
+The final Appendix H gives some examples of multiple answers generated by StochA.
 Appendix A Architecture detail
@@
 train inference Swi+FT StochA: p Swi+FT: $\alpha$ sparsity (%)
-baselines SILU SILU 00000.0002
-RELU RELU 94.8
-Swi+FT SILU RELU 0.05 79.9
+baselines SILU SILU - 00000.0002
+RELU RELU - 94.8
+Swi+FT SILU RELU - 0.05 79.9
 StochA [S|R]-S+ RELU 0.3 0.05 88.5
@@
 [S|R]-S+ RELU 0.7 0.05 84.6
-Table 5: We report the rate of the 0-valued activation after the SILU (not sparse) and RELU activation, when training with the LM1.5B model with the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$.
-Pre-activation inputs distribution dead neurons in the $W_{1}$ layer
+Table 5: We report the rate of the 0-valued activation after the SILU (not sparse) and RELU activation, when training with the LM1.5B model with the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values of $p$.
+Pre-activation inputs distribution Presence of dead neurons in the $W_{1}$ layer
 Refer to caption Refer to caption
@@
 Right: fraction of near-zero rows of the $W_{1}$ matrix, indicating useless neurons for two types of activations.
-In Table 5, we report the sparsity rates resulting from the gating in the FFN layer, when using the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$.
+In Table 5, we report the sparsity ratios resulting from the gating in the FFN layer, when using the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$.
 In all cases, except for the SILU baseline, we use RELU at test time.
@@
 For $x<0$, the probability $P(X\leq x)$ is lower for SILU than for both Swi+FT and when we combine StochA with Swi+FT for different values of $p$.
-This shows that combining StochA with Swi+FT $P(X\leq x)$ enables us to control how much mass we assign to negative values of $x$ depending on $p$.
-The sudden jump in the CDF for RELU indicates that there are many activations that are exactly 0 on input of RELU.
-When investigating the rows of linear layer $W_{1}$ (Eq 5) in Figure 7 (right), it turns out a large fraction of these are near 0 (i.e.
-their norm is less than 1/1000 the average row norm of the matrix).
-This means that these rows are unused: they are “dead neurons”.
-Fortunately, with our Swi+FT approach, that uses SILU at pre-training, these dead neurons disappear.
-This explains why our approach can obtain significantly better results with RELU than the vanilla training of the model with the RELU activation employed from scratch.
-Appendix C Benchmarks
+This shows that combining StochA with Swi+FT enables us to control how much mass we assign to negative values of $x$ depending on $p$.
+Appendix C Diversity of generations ablation experimental setup
+We use the LM3B models trained with the StochA activation [S|R]-S+ with $p$ in $\{0.3,0.5,0.7,0.85\}$.
+At inference time, the probability $p$ is set to be the same as the one during training.
+We generate twenty answers for the NQ and TQA generation tasks, compute the Ff1 and Type-token ratio (TTR) metrics of $n$ generated answers (samples) and average over all datapoints for a given number of samples $n$ in $\{2,3,4,5,10,20\}$.
+In Figure 6 (b) we report the type-token ratio metric (Johnson, 1944), a simple and widely used metric in linguistics and natural language processing to measure lexical diversity in a text.
+In this context, a type consists of a unique word in the text and a token consists of any word occurrence in the text (including repetitions).
+The Type-Token Ratio is then TTR=Number of Types / Number of Tokens.
+Appendix D Multiple forward passes per generated token
+If we use the StochA activation function, it is equivalent to doing Monte Carlo dropout (Gal & Ghahramani, 2016) on the negative side of the input after activation: it is set to zero with probability 1p1-p.
+For this reason, at test time, if we run multiple forward passes, this naturally creates a random output after activation which is either zero or SILU on the negative side of the inputs after activation.
+In this setup, at generation time, after performing the multiple forward passes, we select the next token according to the following strategies:
+Max probability: we select the token that has the highest probability.
+Majority vote: we select the next token based on the token which appears the most often across the multiple passes.
+Results.
+We use LM3B models trained with the [S|R]-S+ activations with $p$ in $\left\{0.3,0.5,0.7,0.85\right\}$.
+We vary the number of forward passes in $\left\{5,10,15,20,30\right\}$.
+We report the average performance of the generation tasks nq, tqa, mbpp and human eval plus and greedy decoding.
+Our findings indicate that even though we observe more diversity at the token level, the majority vote and maximum probability strategies do not perform better than the baseline, where we have a single forward pass as we increase the number of forward passes.
+Appendix E Continuous Pretraining LM1.5B models
+We can leverage the StochA and Swi+FT methods for continuous pretraining (CPT) language models, besides using these activations to train LMs from scratch.
+In this experiment, we start from the LM1.5B models pretrained with the SILU activation.
+During CPT, we use i) Swi+FT ii) StochA with $p=0.3$ or $p=0.5$ or iii) StochA and Swi+FT with $p=0.5$ and $\alpha=0.1$.
+We use the same number of steps used in the pretraining experiments: 45,629 steps.
+We compare the training losses against the baseline model that keeps SILU during CPT.
+pretrain CPT inference StochA: p sparsity (%) train loss
+baseline SILU SILU SILU - 0.0002 2.076
+Swi+FT SILU RELU RELU - 90.96 2.086
+StochA SILU StochA RELU 0.5 41.02 2.095
+SILU StochA RELU 0.3 57.93 2.095
+StochA +Swi+FT SILU StochA +Swi+FT RELU 0.5 81.61 2.096
+Table 6: We report the rate of the 0-valued activations and training losses when doing CPT with the LM1.5B model with the following activations: RELU, SILU, Swi+FT,StochA and when using StochA with Swi+FT for varying values $p$.
+Appendix F StochA and Swi+FT for (TANH, RELU) activations
+Our StochA and Swi+FT methods can be generalized, starting with any pair of non-sparse, sparse activations, in the following way:
+  1. 1.
+    StochA: the non-sparse (activation which takes non-zero values on the negative side) and sparse activation pair can be used as a stochastic activation function that takes the form of either of the activations on the positive side, and, on the negative side, it either takes the non-zero value of the non-sparse variant with probability $p$ or is set to zero with probability 1p1-p.
+  2. 2.
+    Swi+FT: we start with an LM with a non-sparse activation and train it for $(1-\alpha)\%$ steps. Successively, we finetune the LM with the sparse activation for $\alpha\%$ of the final steps.
+To illustrate how the StochA and Swi+FT methods are generalizable, we use the pair of (TANH, RELU) activations.
+We pretrain a LM1.5B with TANH as the non-sparse activation and RELU as the sparse one for 1) and 2).
+We compare these versus the losses to the corresponding baselines: using only the TANH or only the RELU activations, respectively.
+We denote the activation for 1) as Tanh+FT with parameter $\alpha$ and for 2), with [T|R]-T+ with parameter $p$.
+We observe that there is barely any gap between the train loss of the non-sparse (TANH, bolded) and the sparse (RELU, in red) activations, hence, there is no trade-off between sparsity and performance, in contrast to the (SILU, RELU) case.
+Training activation LM1.5B
+$x<0$ $x>0$ $p$ train val val
+Activation pp 1p1-p Swi+FT RELU StochA
+TANH TANH TANH - 2.133 2.155
+RELU RELU RELU - 2.140 2.162
+Tanh+FT TANH TANH - 2.193 2.224
+[T|R]-T+ TANH RELU TANH 0.3 2.368 10.803 2.421
+[T|R]-T+ TANH RELU TANH 0.5 2.314 10.997 2.359
+Table 7: Losses: train is computed over the last 500 steps of the training loss of LM1.5B, val is measured after training on a different set of text and code using the RELU activation or StochA, i.e., the same activation used at train time (possibly deterministic).
+If Swi+FT is enabled, we switch to RELU for the last $10\%$ steps.
+Appendix G Benchmarks
 Code generation
@@
   • MBPP (Austin et al., 2021) is designed to evaluate the code generation abilities of AI models, particularly for Python programming tasks.
+Benchmark Metric Few Shot Type
+hellaswag acc_char 0 choice
+winogrande acc_char 0 choice
+arc_easy acc_char 0 choice
+arc_challenge acc_char 0 choice
+piqa acc_char 0 choice
+obqa acc_char 0 choice
+race.middle acc_char 0 choice
+race.high acc_char 0 choice
+human_eval_plus pass@1 0 generation
+mbpp compiles@1 3 generation
+tqa f1 5 generation
+nq f1 5 generation
+Table 8: Summary of benchmarks used for evaluation. We few-shot prompt the LM1.5B and LM3B models. The metric depends on the benchmark as well as the average performance.
 Common sense and general reasoning
@@
 The benchmark is designed to test multi-step mathematical reasoning capabilities in language models.
-Benchmark Metric Few Shot Type
-hellaswag acc_char 0 choice
-winogrande acc_char 0 choice
-arc_easy acc_char 0 choice
-arc_challenge acc_char 0 choice
-piqa acc_char 0 choice
-obqa acc_char 0 choice
-race.middle acc_char 0 choice
-race.high acc_char 0 choice
-human_eval_plus pass@1 0 generation
-mbpp compiles@1 3 generation
-tqa f1 5 generation
-nq f1 5 generation
-Table 6: Summary of benchmarks used for evaluation. We few-shot prompt the LM1.5B and LM3B models. The metric depends on the benchmark as well as the average performance.
-Appendix D Examples of various sequences generated with StochA
-Using a pre-trained LM3B model with StochA, we can use the StochA activation at test time to generate multiple predictions by leveraging the randomness from the activation function.
-We provide two examples of the generations obtained using questions from the TQA benchmark (Joshi et al., 2017):
-question and generated number of NLL F1
-accepted answers answers times score score
+Appendix H Examples of various sequences generated with StochA
+We use a pre-trained LM3B model with StochA, we can use the StochA activation at test time to generate multiple predictions by leveraging the randomness from the activation function.
+In Table LABEL:tab:exmultigen, we provide two examples of the generations obtained using questions from the TQA benchmark (Joshi et al., 2017):
+Table 9: Example generations.
+For each question, we indicate the ground-truth answers (from the dataset).
+We generate 20 answers per question with [S|R]-S+, pp=0.7.
+We list the de-duplicated answers, with the NLL score (used to sort the results) and the F1 score (used to evaluate the result, it is computed as the intersection of bags-of-words).
 When did ibuprofen become available over the counter? A: 1983 A: 1984 1988 4 1.002 0
@@
 Nasion 3 1.224 0
-Table 7: Example generations.
-For each question, we indicate the ground-truth answers (from the dataset).
-We generate 20 answers per question with [S|R]-S+, pp=0.7.
-We list the de-duplicated answers, with the NLL score (used to sort the results) and the F1 score (used to evaluate the result, it is computed as the intersection of bags-of-words).